OpenStudy (anonymous):

Given the function f described by f(y)=4y^2+9y-3 Find the function values

OpenStudy (anonymous):

0=16Y+9Y-3 0=25Y-3 3+ +3 ------------ 3=25Y 3/25=25Y/25 Y=25/3

OpenStudy (anonymous):

I THINK THIS IS THE ANSWER!!!!!!!!!!!

OpenStudy (anonymous):

are you saying the answer is 25/3